[figure: 9.7 cm]

THE STRUCTURE AND MAGNETIC PROPERTIES OF Fe-Si

POWDER PREPARED BY BALL MILLING

J. Füzer^{a} P. Kollár ^{a} P. Petrovič^{a} I. Tóth^{b} V. Kavečanský^{c} J. Kováč^{c}

and T. Švec ^{a}

aDept. of Exp. Physics, Fac. of Sci., P.J. Šafárik University Park Angelinum 9, 041 54

Košice, Slovakia bDepartment of Nuclear Physics and Technology, Slovak Tech. Univ., 812 19 Bratislava, Slovakia

Institute of Exp. Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovakia

We have prepared microcrystalline FeSi (6.5wt. % of Si) ribbon by melt spinning method and then milled in vibratory micro-mill up to 800 hours. The structure was detected by SEM and X-ray diffraction. Magnetic moment and coercivity were measured by vibrating sample magnetometer (VSM). Thermomagnetic curves were taken also by VSM. The room temperature Mssbauer transmission spectra were obtained using a conventional spectrometer with ^{57}Co/Rh γ-source. XRD spectra show the presence of two phases. The first phase is iron rich phase and second one is silicon-rich phase. Both phases were detected by measuring of thermomagnetic curves. The coercivity of milled powder at the beginning of milling process steeply increases and riches its maximium at 7.5 kA/m for 500 hours of milling. Than it decreases with milling time very slowly. We suppose that the maximum of the coercivity corresponds to the coercivity of single domain particles of iron rich Fe-Si phase.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Füzer

Address for correspondence : Dept. of Exp. Physics, Fac. of Sci., P.J. Šafárik University Park Angelinum 9, 041 54 Košice, Slovakia For correspondence

Email address : jfuzer@kosice.upjs.sk